OpenStudy (anonymous):

What is the range of y = 5x -2 if the domain is {-3, -1, 0, 1, 3}? {-13, -3, 2, 3, 13} {12, 3, 0, 3, 13} {-17, -7, -2, 3, 13}

OpenStudy (anonymous):

substitute x=-3,-1,0,1,3 into the equation @theviolentredhead

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

y=5(-3)-2 y=-15-2 y=?

OpenStudy (anonymous):

solve it @theviolentredhead

OpenStudy (anonymous):

-17 @MARC_